JUST DEW IT CORPORATION 2017 and 2018 Balance Sheets Assets Liabilities and Owners Equity 2017 2018 2017 2018 Current assets Current liabilities Cash Accounts $ 40,000 41,800 21,600 $ 13,280 $14,600 19,200 51,000 $ 63,200 $84,800 Accounts payable Notes payable 7,520 14,240 receivable Inventory 42,400 Total Total $ 54,240 63,400o Long-term debt 32,000 24,000 Owners equity Common stock and paid-in $ 40,000 40,000 surplus 272,600 $233,760 312,600 Total liabilities and owners $320,000 $400,000 Retained earnings 193,760 Net plant and equipment $256,800 315,200 Total Total assets $320,000 400,000 equity
Answer #1

a) In 2017 Current Ratio = Current Assets/ current liabilities = 63200/54240 = 1.17
In 2018 Current Ratio = Current Assets/ current liabilities = 84800/63400 = 1.34

b) In 2017 Quick Ratio = (Current Assets-Inventories)/ current liabilities = (63200-42400)/54240 = 0.38
In 2018 Quick Ratio = (Current Assets-Inventories)/ current liabilities = (84800-51000)/63400 = 0.53

c) In 2017 Cash ratio = Cash/ Current Liabilities = 13280/54240 = 0.24
In 2018 Cash ratio = Cash/ Current Liabilities = 14600/63400 = 0.23

d) In 2017 NWC to Total Asset ratio = ( Current Asset - Current Liabilities)/Total Assets = (63200-54240)/320,000 = 0.03
In 2018 NWC to Total Asset ratio = ( Current Asset - Current Liabilities)/Total Assets = (84800-63400)/400,000 = 0.05